I have to write an equation relating the number of cups of brown sugar (s) to the number of cups of flour (f) in a recipe that calls for 2 1/4 cups of all purpose flour and 3/4 cup brown sugar.
Answer by AnlytcPhil(1116)

I have to write an equation relating the number of cups of brown sugar (s) to
the number of cups of flour (f) in a recipe that calls for 2 1/4 cups of all
purpose flour and 3/4 cup brown sugar.

The number of cups of flour is directly proportional to the number of cups of
brown sugar, so we first write:

f = ks

when s = 3/4, f = 2 1/4 = 9/4

Substitute these values in f = ks 

9/4 = k(3/4)

Multiply both sides by 4

9 = 3k

3 = k

Rewrite f = ks using 3 for k

f = 3s

That's the answer

Edwin
AnlytcPhil@aol.com